A furniture shop refinishes tables. Employees use two methods to refinish tables. Method I takes
0.5
hours and the material costs
$9
.
Method II takes
2.0 hours, and the material costs
$7
.
Next week, they plan to spend
99 hours in labor and
$564 in material for refinishing tables. How many tables should they plan to refinish with each method? (Round to two decimal places if necessary

Answer by jorel1380(3719)

Let the two methods be x and y. Then:
.5x+2y=99
9x+7y=564
So:
9x+36y=1782
9x+7y=564
29y=1218
y=42
x=30
They should refinish 30 tables using the first method, and 42 tables using the second. ☺☺☺☺